OpenStudy (anonymous):

How can the properties of rational exponents be applied to simplify expressions with radicals or rational exponents?

ganeshie8 (ganeshie8):

see if you can go through that link and come up with some explanation on how to apply the properties

ganeshie8 (ganeshie8):

this link has nice explanaiton http://hotmath.com/hotmath_help/topics/properties-of-exponents.html

ganeshie8 (ganeshie8):

start by something like belpw : the rational exponent properties can be used to simplify radical expressions by writing the radical as exponent and simplifying the exponent, for example : \[\huge \sqrt[3]{a^6}\] can be written as : \[\huge \left(a^6\right)^{\frac{1}{3}} \] which simplifies to : \[\huge a^{\frac{6}{3}} = a^2 \]

ganeshie8 (ganeshie8):

keep going... write 2-3 more paragraphs and finish it off

OpenStudy (anonymous):

ty again